A jar contains 5 red marbles, 5 white marbles, and 6 blue marbles. Two marbles are randomly selected from the jar with replacement. Find the following probabilities. Leave your answer as a reduced fraction.




1
Expert's answer
2022-02-15T12:48:57-0500

Total number of marbles is 5+5+6=16 marbles

Let the events, "R,W,B" representing red, white and blue marble selected respectively.

Probability of choosing a red marble is "{5\over16}"

Probability of choosing a white marble is "{5\over16}"  

Probability of choosing a blue marble is "{6\over 16}"


"a)"

The probability that a red marble and then a blue marble is randomly selected is "{5\over16}\times{6\over16}={30\over 256}={15\over`128}"     


"b)"

The probability that a white marble and then a red marble is randomly selected is, "{5\over16}\times{5\over16}={25\over256}"